A basketball player scored 25 times during one game. He scored a total of 48 points, two for each two-point shot and one for each free throw. How many two-point shots did he make? How many free throws?
___ two point shots
____ free throws

Answer by josgarithmetic(39618)

let x mean how many two-point shots
let y mean how many one-point free throw

x+y=25 accounts only for how many scorings this player made
x*2+y*1=48 accounts for number of points scored

System to solve:
2x+y=48
x+y=25

Subtract second from first,
x+0=48-25
x=23
....y=2